# Dynamical Low-Rank Compression of Neural Networks with Robustness under Adversarial Attacks

### v1

@@ Abstract
Deployment of neural networks on resource-constrained devices demands models that are both compact and robust to adversarial inputs. However, compression and adversarial robustness often conflict. In this work, we introduce a dynamical low-rank training scheme enhanced with a novel spectral regularizer that controls the condition number of the low-rank core in each layer. This approach mitigates the sensitivity of compressed models to adversarial perturbations without sacrificing clean accuracy. The method is model- and data-agnostic, computationally efficient, and supports rank adaptivity to automatically compress the network at hand. Extensive experiments across standard architectures, datasets, and adversarial attacks show the regularized networks can achieve over 94% compression while recovering or improving adversarial accuracy relative to uncompressed baselines.
@@ 1 Introduction
Deep neural networks have achieved state-of-the-art performance across a wide range of tasks in computer vision and data processing. However, their success comes at the cost of substantial computational and memory demands which hinder deployment in resource-constrained environments. While significant progress has been made in scaling up models through data centers and specialized hardware, a complementary and equally important challenge lies in the opposite direction: deploying accurate and robust models on low-power platforms such as unmanned aerial vehicles (UAVs) or surveillance sensors. These platforms often operate in remote locations with limited power and compute resources, and are expected to function autonomously over extended periods without human intervention.
@@
Recent work has shown that these three objectives are inherently at odds. Compression via low-rank steffen2022low or sparsity techniques guo2016dynamic often leads to reduced accuracy. Techniques to improve adversarial robustness—such as data augmentation lee2023generativeadversarialtrainerdefense or regularization-based defenses zhang2019theoreticallyprincipledtradeoffrobustness —frequently degrade clean accuracy. Moreover, it has been observed that compressed networks can exhibit increased sensitivity to adversarial attacks savostianova2023robustlowranktrainingapproximate . Finally, many methods to increase adversarial robustness of the model impose additional computational burdens during training 10302664 ; 9442932 or inference pmlr-v70-cisse17a ; hein2017formal ; liu2019advbnnimprovedadversarialdefense , further complicating deployment on constrained hardware.
@@
  • Low-rank compression framework. We introduce a novel regularizer and training framework that yields low-rank compressed neural networks, achieving a more than 10×10$\times$10 × reduction in both memory footprint and compute cost, while maintaining clean accuracy and adversarial robustness on par with full-rank baselines.
@@
Beyond these core contributions, our approach is model- and data-agnostic, can be integrated seamlessly with existing adversarial defenses, and never requires assembling full-rank weight matrices – the last point guaranteeing a low memory footprint during training and inference. Moreover, by connecting to dynamical low-rank integration schemes and enabling convergence analysis via gradient flow, we offer new theoretical and algorithmic insights. Finally, the use of interpretable spectral metrics enhances the trustworthiness and analyzability of the compressed models.
@@ 2 Related work
In contrast, low-rank fine-tuning methods like LoRA (hu2021lora, ) inject trainable low-rank updates into a frozen pre-trained model, enabling efficient adaptation with few parameters. Extensions such as GeoLoRA schotthöfer2024geolorageometricintegrationparameter , AdaLoRA (zhang2023adalora, ), DyLoRA (valipour2023dylora, ), and DoRA (Mao2024DoRAEP, ) incorporate rank adaptation or structured updates, improving performance over static rank baselines. However, these fine-tuning methods do not reduce the cost of the full training and inference, thus are not applicable to address the need of promoting computational efficiency.
@@
Many hard approaches strongly enforce orthogonality/well-conditioned constraints by training on a chosen manifold using Riemannian optimization methods li2020efficientriemannianoptimizationstiefel ; Projection_like_Retractions ; savostianova2023robustlowranktrainingapproximate . A hard approach built for low-rank training is given in savostianova2023robustlowranktrainingapproximate ; this method clamps the extremes of the spectrum to improve the condition number during the coefficient update. The clamping gives a hard estimate on the range of the spectrum which enables a direct integration of the low-rank equations of motion with standard learning rates. However, this method requires a careful selection of the rank $r$, which is viewed as a hyperparameter in savostianova2023robustlowranktrainingapproximate . If $r$ is chosen incorrectly, the clamping of the spectrum, a hard-thresholding technique, acts as a strong regularizer which could affect the validation metrics of the network.
@@
Our method falls neatly into a soft approach and our proposed regularizer can be seen as an extension of the soft orthogonality (SO) regularizer xie2017all to well-conditioned matrices in the low-rank setting. As noted in bansal2018can , the SO regularizer only works well when the input matrix is of size $m\times n$ where $m\leq n$. However, in our case we avoid this issue since the low-rank coefficient matrix $S$ is always square; an extension to convolutional layers is discussed in (4.1). In the context of low-rank training, the soft approach enables rank-adaptivity of the method.
We consider a neural network $f$ as a concatenation of $L$ layers $z^{\ell+1}=\sigma^{\ell}(W^{\ell}z^{\ell})$ with matrix valued [^1] We provide an extension to tensor-valued layers, e.g. in CNNs, in Section 4.1 parameters $W^{\ell}\in\mathbb{R}^{n\times n}$, layer input $z^{\ell}\in\mathbb{R}^{b\times n}$ and element-wise nonlinear activation $\sigma^{\ell}$. We do not consider biases, but a model with biases can always be formulated without them by folding them into the weights and creating a dimension that is always one. The data $X$ constitutes the input to the first layer, i.e. $z^{0}=X$. We assume that the layer activations $\sigma^{\ell}$ are Lipschitz continuous, which is the case for all popular activations savostianova2023robustlowranktrainingapproximate . The network is trained on a loss function $\mathcal{L}$ which we assume to be locally bounded and Lipschitz continuous gradient.
Low-rank Compression: The compression the network for training and inference is typically facilitated by approximating the layer weight matrices by a low-rank factorization $W^{\ell}=U^{\ell}S^{\ell}V^{\ell,\top}$ with $U^{\ell},V^{\ell}\in\mathbb{R}^{n\times r}$ and $S^{\ell}\in\mathbb{R}^{r\times r}$, where $r$ is the rank of the factorization. In this work, we generally assume that $U^{\ell},V^{\ell}$ are orthonormal matrices at all times during training and inference. This assumption deviates from standard low-rank training approaches hu2021lora , however recent literature provides methods that are able to fulfill this assumption approximately zhang2023adalora and even exactly steffen2022low ; schotthöfer2024geolorageometricintegrationparameter . If $r\ll n$, the low-rank factorization with computational cost of $\mathcal{O}(2nr+r^{2})$ is computationally more efficient than the standard matrix format $W$ with computational cost of $\mathcal{O}(n^{2})$. Throughout this work, we call a network in the standard format a "baseline" network.
Adversarial robustness: The adversarial robustness of a neural network $f$, a widely used trustworthiness metric, can be measured by its relative sensitivity $\mathcal{S}$ to small perturbations $\delta$, e.g., noise, of the input data $X$ 9726211 ; 10.1145/3274895.3274904 , i.e., $\mathcal{S}(f,X,\delta):=\tfrac{||f(X+\delta)-f(X)||}{||f(X)||}\tfrac{||X||}{||\delta||}$.
In this work, we consider the sensitivity in the Euclidean ($\ell^{2}$) norm, i.e., $||\cdot||=||\cdot||_{2}$. For neural networks consisting of layers with Lipschitz continuous activation functions $\sigma^{\ell}$, $\mathcal{S}$ can be bounded savostianova2023robustlowranktrainingapproximate by the product
$\vspace{-0.1cm}\mathcal{S}(f,X,\delta)\leq\big{(}\textstyle\prod_{\ell=1}^{L}\kappa(W^{\ell})\big{)}\big{(}\prod_{\ell=1}^{L}\kappa(\sigma^{\ell})\big{)}\vspace{-0.1cm}$ (1)
Figure 1: The singular values $\sigma_{i}(W)$ of sequential layer 7 in VGG16 for baseline training, dynamical low-rank training (DLRT), and regularized DLRT with our condition number regularizer $\mathcal{R}$ with $\beta=0.075$. The matrix $W$ is formed as the first-mode unfolding of the convolutional tensor. Conditioning of the regularized low-rank layer is significantly improved compared to the non-regularized low-rank and baseline layer.2
where $\kappa(W^{\ell}):=\left\lVert W^{\ell}\right\rVert\left\lVert W^{\ell,{\dagger}}\right\rVert$ is the condition number of the matrix of layer $\ell$, $W^{\dagger}$ is the pseudo-inverse of a matrix $W$, and $\kappa(\sigma^{\ell})$ is the condition number of the layer activation function $\sigma^{\ell}$. The condition number of the element-wise non-linear activation functions $\sigma^{\ell}$ can be computed with the standard definitions (see trefethen1997numerical and savostianova2023robustlowranktrainingapproximate for condition numbers of several popular activation functions). Equation 1 allows us to consider each layer individually, thus we drop the superscript $\ell$ for brevity of exposition.
The sensitivity of a low-rank factorized network can be readily deducted from Equation 1 by leveraging orthonormality of $U,V$: $\kappa(USV^{\top})=\left\lVert USV^{\top}\right\rVert\left\lVert VS^{\dagger}U^{\top}\right\rVert=\left\lVert S\right\rVert\left\lVert S^{\dagger}\right\rVert=\kappa(S)$. This fact allows us to only consider the $r\times r$ coefficient matrix $S$ to control the networks sensitivity. The condition number $\kappa(S)$ can be determined via a singular value decomposition (SVD) of $S$, which is computationally feasible when $r\ll n$.
Adversarial robustness-aware low-rank training: Enhancing the adversarial robustness of the network during low-rank training thus boils down to control the conditioning of $S$, which is a non-trivial task. In fact, the dynamics of the singular spectrum of $S$ of adaptive low-rank training schemes as DLRT steffen2022low exhibit ill conditioned behavior, even if $S$ is always full rank. In Figure 1, we observe that the singular values a rank 64 factorization of a network layer compressed with DLRT range from $\sigma_{r=1}=2.7785$ to $\sigma_{r=64}=0.8210$ yielding a condition number of $\kappa(S)=3.3844$. In comparison, the baseline network has singular values ranging from $\sigma_{r=1}=1.8627$ to $\sigma_{r=128}=0.9445$ yielding a lower condition number of $\kappa(S)=1.9722$. As a result, a $\ell^{2}$-FGSM attack with strength $\epsilon=0.3$, reduces the accuracy of the baseline network to $54.96\%$, while the low-rank network’s accuracy drops to $43.39\%$, see Table 2.
@@
Thus, we design a computationally efficient regularizer $\mathcal{R}$ to control and decrease the condition number of each network layer during training. The regularizer $\mathcal{R}$ only acts on the small $r\times r$ coefficient matrices $S$ of each layer and thus has a minimal memory and compute overhead over low-rank training. The regularizer is differentiable almost everywhere and compatible with automatic differentiation tools. Additionally, ${\mathcal{R}}$ has a closed form derivative that enables an efficient and scalable implementation of ${\nabla}{\mathcal{R}}$. Furthermore, ${\mathcal{R}}$ compatible with any rank-adaptive low-rank training scheme that ensures orthogonality of $U,V$, e.g., zhang2023adalora ; schotthöfer2024federateddynamicallowranktraining ; schotthöfer2024geolorageometricintegrationparameter ; savostianova2023robustlowranktrainingapproximate .
@@
The regularizer ${\mathcal{R}}$ can be viewed as an extension of the soft orthogonal regularizer xie2017all ; bansal2018can where we penalize the distance of $S^{\top}\!S$ to the well-conditioned matrix $\alpha_{S}^{2}I$. Here $\alpha_{S}$ is chosen such that $\|S\|=\|\alpha_{S}I\|$. Moreover, ${\mathcal{R}}$ is also a scaled standard deviation of the set $\{\sigma_{i}(S)^{2}\}_{i=1}^{r}$, that is,
$\frac{1}{r}{\mathcal{R}}(S)^{2}=\frac{1}{r}\sum_{i=1}^{r}(\sigma_{i}(S)^{2})^{2}-\Big{(}\frac{1}{r}\sum_{i=1}^{r}\sigma_{i}(S)^{2}\Big{)}^{2}.$ (3)
See Appendix B for the proof. Therefore, ${\mathcal{R}}$ is a unitarily invariant regularizer; namely, ${\mathcal{R}}(USV^{\top})={\mathcal{R}}(S)$ for orthogonal $U,V$. These two forms of ${\mathcal{R}}$ are useful in the properties shown below.
@@
See Appendix B for the proof. The gradient computation consists only of $r\times r$ matrix multiplications and an Frobenius norm evaluation, and is thus computationally efficient for $r\ll m$. Further, its closed form enables a straight-forward integration into existing optimizers as Adam or SGD applied to $S$.
Table 1: VGG16 on UCM data. Comparison of adversarial robustness of LoRA and DLRT trained networks under the $\ell^{2}$-FGSM attack with similar compression rate. Orthogonality of $U,V$ increases adversarial performance significantly.
@@
We have the condition number bound
$\kappa(S)\leq\exp\bigg{(}\frac{1}{\sqrt{2}\sigma_{r}(S)^{2}}{\mathcal{R}}(S)\bigg{)}$ (4)
for any $S\in\mathbb{R}^{r\times r}$.
See Appendix B for the proof. Thus, if $\sigma_{r}(S)$ is not too small, we can use ${\mathcal{R}}(S)$ as a good measure for the conditioning of $S$. Note that the singular value truncation of Algorithm 1 ensures that $\sigma_{r}(S)$ is always sufficiently large. Figure 2a) and b) show the dynamics of ${\mathcal{R}}(S(t))$ and $\kappa(S(t))$ during regularized training using Algorithm 1, and we see that $\kappa(S(t))$ decays as ${\mathcal{R}}(S(t))$ decays, validating 2.
Remark 1.
When $U,V$ are not orthonormal, e.g. in simultaneous gradient descent training (LoRA), we have ${\mathcal{R}}(USV^{\top})\not={\mathcal{R}}(S)$. $USV^{\top}$ are zero-valued thus the bound of Equation 4 is not useful. Table 1 shows that the clean accuracy and adversarial accuracy of regularized LoRA is significantly lower than baseline training or regularized training with orthonormal $U,V$.
Figure 2: UCM Dataset, $\kappa(S(t))$ and $\mathcal{R}(S(t))$ of layer 4 of VGG16 for different regularizations strengths $\beta$. Each line is the median of 5 training runs. Higher $\beta$ values lead to faster reduction of the layer condition $\kappa(S)$, which quickly approaches its minimum value $1$, and faster decay of $\mathcal{R}$. Unregularized training, i.e. $\beta=0$ leads to $\kappa(S)>1e3$ after few iterations.
We now study the stability of the regularizer when applied to a least squares regression problem, i.e., given a fixed $W\in\mathbb{R}^{r\times r}$ we seek to minimize $\mathcal{J}(S):=\beta{\mathcal{R}}(S)+\frac{1}{2}\|S-W\|^{2}$ over $S\in\mathbb{R}^{r\times r}$.
Proposition 3.
For a fixed $W\in\mathbb{R}^{r\times r}$, consider the dynamical system $\dot{S}(t)+\beta{\nabla}{\mathcal{R}}(S(t))+S(t)=W$. Then for any $t\geq 0$ we have the long-time stability estimate
$\tfrac{1}{2}\|S(t)-W\|^{2}+2\beta\int_{0}^{t}e^{\tau-t}{\mathcal{R}}(S(\tau))\,\mathrm{d}\tau\leq\tfrac{1}{2}e^{-t}\|S(0)-W\|^{2}+2(1-e^{-t})\beta(1+2\beta)\|W\|^{2}.$ (5)
See Appendix B for the proof. We note that unlike standard ridge and lasso regularizations methods, ${\mathcal{R}}$ lacks convexity; thus long-time stability of the regularized dynamics is not obvious. However, ${\nabla}{\mathcal{R}}$ possesses monotonicity properties that we leverage to show in (5) that the growth in ${\mathcal{L}}$ in the regularized problem only depends on $\beta$, $W$, and the initial loss. Moreover, for large $t$, the change in the final loss by the regularizer only depends on $\beta$ and the true solution $W$ and not the specific path $S(t)$. While training on the non-convex loss will not provide the same theoretical properties as the least-square loss, 3 and the experiments in Figure 2 give confidence that adding our regularizer does not yield a relatively large change in the loss decay rate over moderate training regimes. Particularly, we observe empirically in Figure 2 that the condition number $\kappa(S)$ of decreases alongside the regularizer value $\mathcal{R}$ during training.

### v2

@@ Abstract
Deployment of neural networks on resource-constrained devices demands models that are both compact and robust to adversarial inputs. However, compression and adversarial robustness often conflict. In this work, we introduce a dynamical low-rank training scheme enhanced with a novel spectral regularizer that controls the condition number of the low-rank core in each layer. This approach mitigates the sensitivity of compressed models to adversarial perturbations without sacrificing accuracy on clean data. The method is model- and data-agnostic, computationally efficient, and supports rank adaptivity to automatically compress the network at hand. Extensive experiments across standard architectures, datasets, and adversarial attacks show the regularized networks can achieve over 94% compression while recovering or improving adversarial accuracy relative to uncompressed baselines.
@@ 1 Introduction
Deep neural networks have achieved state-of-the-art performance across a wide range of tasks in computer vision and data processing. However, their success comes at a cost of substantial computational and memory demands, which hinders deployment in resource-constrained environments. While significant progress has been made in scaling up models through data centers and specialized hardware, a complementary and equally important challenge lies in the opposite direction: deploying accurate and robust models on low-power platforms such as unmanned aerial vehicles (UAVs) or surveillance sensors. These platforms often operate in remote locations with limited power and compute resources, and are expected to function autonomously over extended periods without human intervention.
@@
Recent work has shown that these three objectives are inherently at odds. Compression via low-rank steffen2022low or sparsity techniques guo2016dynamic often leads to reduced accuracy. Techniques to improve adversarial robustness—such as data augmentation lee2023generativeadversarialtrainerdefense or regularization-based defenses zhang2019theoreticallyprincipledtradeoffrobustness —frequently degrade clean accuracy. Moreover, it has been observed that low-rank compressed networks can exhibit increased sensitivity to adversarial attacks savostianova2023robustlowranktrainingapproximate . Finally, many methods to increase adversarial robustness of the model impose additional computational burdens during training 10302664 ; 9442932 or inference pmlr-v70-cisse17a ; hein2017formal ; liu2019advbnnimprovedadversarialdefense , further complicating deployment on constrained hardware.
@@
  • Low-rank compression framework. We introduce a novel regularization and incorporation method to modify a class of low-rank training frameworks that yields low-rank compressed neural networks, achieving a more than $10\times$ reduction in both memory footprint and compute cost, while maintaining clean accuracy and adversarial robustness on par with full-rank baselines.
@@
Beyond these core contributions, our approach is model- and data-agnostic, can be integrated seamlessly with existing adversarial defenses, e.g., adversarial training goodfellow2014explaining , and never requires assembling full-rank weight matrices – the last point guaranteeing a low memory footprint during training and inference. Moreover, by connecting to dynamical low-rank integration schemes and enabling convergence analysis via gradient flow, we offer new theoretical and algorithmic insights. Finally, the use of interpretable spectral metrics enhances the trustworthiness and analyzability of the compressed models.
We consider a neural network $f$ as a concatenation of $L$ layers $z^{\ell+1}=\sigma^{\ell}(W^{\ell}z^{\ell})$ with matrix valued [^1] We provide an extension to tensor-valued layers, e.g. in CNNs, in Section 5.1 parameters $W^{\ell}\in\mathbb{R}^{n\times n}$, layer input $z^{\ell}\in\mathbb{R}^{n\times b}$ and element-wise nonlinear activation $\sigma^{\ell}$. For simplicity of notation, we do not consider biases, but they are included in the numerical experiments in Section˜6. The data $X$ constitutes the input to the first layer, i.e. $z^{0}=X$. We assume that the layer activations $\sigma^{\ell}$ are Lipschitz continuous, which is the case for all popular activations savostianova2023robustlowranktrainingapproximate . The network is trained on a loss function $\mathcal{L}$ which we assume to be locally bounded with a Lipschitz continuous gradient. Throughout this work, we call a network in the standard format a "baseline" network.
Low-rank Compression: The compression the network for training and inference is typically facilitated by approximating the layer weight matrices by a low-rank factorization $W^{\ell}=U^{\ell}S^{\ell}V^{\ell,\top}$ with $U^{\ell},V^{\ell}\in\mathbb{R}^{n\times r}$ and $S^{\ell}\in\mathbb{R}^{r\times r}$, where $r$ is the rank of the factorization. In this work, we generally assume that $U^{\ell},V^{\ell}$ are orthonormal matrices at all times during training and inference. This assumption deviates from standard low-rank training approaches hu2021lora , however recent literature provides methods that are able to fulfill this assumption approximately zhang2023adalora and even exactly steffen2022low ; schotthöfer2024geolorageometricintegrationparameter . If $r\ll n$, the low-rank factorization with a storage and matrix-vector computation cost cost of $\mathcal{O}(2nr+r^{2})$ is computationally more efficient than the standard matrix format $W$ with a computational cost of $\mathcal{O}(n^{2})$.
[…]
$\vskip-2.84544pt\mathcal{S}(f,X,\delta)\leq\big(\textstyle\prod_{\ell=1}^{L}\kappa(W^{\ell})\big)\big(\prod_{\ell=1}^{L}\kappa(\sigma^{\ell})\big)\vskip-2.84544pt$ (1)
Figure 1: The singular values $\varsigma_{i}(W)$ of sequential layer 7 in VGG16 for baseline training, unregularized dynamical low-rank training (DLRT), and RobustDLRT with our condition number regularizer $\mathcal{R}$ with $\beta=0.075$. The matrix $W$ is formed as the first-mode unfolding of the convolutional tensor. Conditioning of the regularized low-rank layer is significantly improved compared to the non-regularized low-rank and baseline layer.2
where $\kappa(W):=\left\lVert W\right\rVert\left\lVert W^{{\dagger}}\right\rVert$ is the condition number of a matrix $W$, $W^{\dagger}$ is the pseudo-inverse of $W$, and $\kappa(\sigma)$ is the condition number of the layer activation function $\sigma$. The condition number of the element-wise non-linear activation functions $\sigma^{\ell}$ can be computed with the standard definitions (see trefethen1997numerical and savostianova2023robustlowranktrainingapproximate for condition numbers of several popular activation functions). Equation˜1 allows us to consider each layer individually, thus we drop the superscript $\ell$ for brevity of exposition.
The sensitivity of a low-rank factorized network can be readily deducted from Equation˜1 by leveraging orthonormality of $U$ and $V$, i.e., $\kappa(USV^{\top})=\kappa(S)$. This fact allows us to only consider the $r\times r$ coefficient matrix $S$ to control the sensitivity of the network. The condition number $\kappa(S)$ can be determined via a singular value decomposition (SVD) of $S$, which is computationally feasible when $r\ll n$.
Adversarial robustness-aware low-rank training: Enhancing the adversarial robustness of the network during low-rank training thus boils down to controlling the conditioning of $S$, which is a non-trivial task. In fact, the dynamics of the singular spectrum of $S$ of adaptive low-rank training schemes as DLRT steffen2022low become more ill-conditioned than the baseline during training, even if $S$ is always full rank. In Figure˜1, we observe that the singular values $\varsigma$ of a rank 64 factorization of a network layer compressed with DLRT range from $\varsigma_{r=1}=2.7785$ to $\varsigma_{r=64}=0.8210$ yielding a condition number of $\kappa(S)=3.3844$. In comparison, the baseline network has singular values ranging from $\varsigma_{r=1}=1.8627$ to $\varsigma_{r=128}=0.9445$ yielding a lower condition number of $\kappa(S)=1.9722$. As a result, an $\ell^{2}$-FGSM attack with strength $\epsilon=0.3$, reduces the accuracy of the baseline network to $54.96\%$, while the low-rank network’s accuracy drops to $43.39\%$, see Table˜2.
@@ 3 Related work
In contrast, low-rank fine-tuning methods like LoRA (hu2021lora, ) inject trainable low-rank updates into a frozen pre-trained model, enabling efficient adaptation with few parameters. Extensions such as GeoLoRA schotthöfer2024geolorageometricintegrationparameter , AdaLoRA (zhang2023adalora, ), DyLoRA (valipour2022dylora, ), and DoRA (Mao2024DoRAEP, ) incorporate rank adaptation or structured updates, improving performance over static rank baselines. However, these fine-tuning methods do not reduce the cost of the full training and inference, thus are not applicable to address the need of promoting computational efficiency.
Pruning is another well studied approach to reduce the number of parameters of a trained neural network jian2022pruningadversariallyrobustneural ; li2022pruningimprovecertifiedrobustness ; NEURIPS2020_e3a72c79 ; pmlr-v188-zhao22a ; chen2022linearitygraftingrelaxedneuron ; jordao2021effect by either sparsifying weight matrices or layer output channels of a network. Typically sparsity pruning is performed after training a fully parametrized neural network and thus only reduces memeory and compute load during inference, while treating training as offline cost.
@@
Many hard approaches strongly enforce orthogonality/well-conditioned constraints by training on a chosen manifold using Riemannian optimization methods li2020efficientriemannianoptimizationstiefel ; Projection_like_Retractions ; savostianova2023robustlowranktrainingapproximate . A hard approach built for low-rank training is given in savostianova2023robustlowranktrainingapproximate ; this method clamps the extremes of the spectrum to improve the condition number during training. The clamping gives a hard estimate on the range of the spectrum which enables a direct integration of the low-rank equations of motion with standard learning rates. However, this method requires a careful selection of the rank $r$, which is viewed as a hyperparameter in savostianova2023robustlowranktrainingapproximate . If $r$ is chosen incorrectly, the clamping of the spectrum, a hard-thresholding technique, acts as a strong regularizer which could affect the validation metrics of the network.
@@
Our regularization method detailed below falls neatly into a soft approach and our proposed regularizer can be seen as an extension of the soft orthogonality (SO) regularizer xie2017all to well-conditioned matrices in the low-rank setting. As noted in bansal2018can , the SO regularizer only works well when the input matrix is of size $m\times n$ wish $m\leq n$. However, in our case we avoid this issue since the regularizer is applied to the square matrix $S$; an extension to convolutional layers is discussed in Section˜5.1. In the context of low-rank training, the soft approach enables rank-adaptivity of the method.
4 Improving conditioning via regularization
@@
We design a computationally efficient regularizer $\mathcal{R}$ to control and decrease the condition number of each network layer during training. The regularizer $\mathcal{R}$ only acts on the small $r\times r$ coefficient matrices $S$ of each layer and thus has a minimal memory and compute overhead over low-rank training. The regularizer is differentiable almost everywhere and compatible with automatic differentiation tools. Additionally, $\mathcal{R}$} has a closed form derivative that enables an efficient and scalable implementation of $\nabla$$\mathcal{R}$}. Furthermore, $\mathcal{R}$} is compatible with any rank-adaptive low-rank training scheme that ensures orthogonality of $U,V$, e.g., zhang2023adalora ; schotthöfer2024federateddynamicallowranktraining ; schotthöfer2024geolorageometricintegrationparameter ; savostianova2023robustlowranktrainingapproximate .
@@
The regularizer $\mathcal{R}$} can be viewed as an extension of the soft orthogonal regularizer xie2017all ; bansal2018can where we penalize the distance of $S^{\top}\!S$ to the well-conditioned matrix $\alpha_{S}^{2}I$. Here $\alpha_{S}$ is chosen such that $\|S\|=\|\alpha_{S}I\|$. Moreover, $\mathcal{R}$} is also a scaled standard deviation of the singular values $\{\varsigma_{i}(S)^{2}\}_{i=1}^{r}$, i.e.,
$\frac{1}{r}{\mathcal{R}}(S)^{2}=\frac{1}{r}\sum_{i=1}^{r}(\varsigma_{i}(S)^{2})^{2}-\Big(\frac{1}{r}\sum_{i=1}^{r}\varsigma_{i}(S)^{2}\Big)^{2}.$ (3)
See Appendix˜C for the proof. Therefore, $\mathcal{R}$} is a unitarily invariant regularizer; namely, ${\mathcal{R}}(USV^{\top})={\mathcal{R}}(S)$ for orthogonal $U,V$. These two forms of $\mathcal{R}$} are useful in the properties shown below.
@@
See Appendix˜C for the proof. The gradient computation consists only of $r\times r$ matrix multiplications and an Frobenius norm evaluation, and is thus computationally efficient for $r\ll m$. Further, its closed form enables a straight-forward integration into existing optimizers as Adam or SGD applied to $S$.
Table 1: VGG16 on UCM data. Comparison of regularized LoRA and DLRT trained networks under the $\ell^{2}$-FGSM attack. Orthogonality of $U,V$ increases adversarial performance significantly.
@@
Proposition 2 (Condition number bound).
For any $S\in\mathbb{R}^{r\times r}$ there holds
$\kappa(S)\leq\exp\bigg(\frac{1}{\sqrt{2}\varsigma_{r}(S)^{2}}{\mathcal{R}}(S)\bigg).$ (4)
See Appendix˜C for the proof. Thus, if $\varsigma_{r}(S)$ is not too small, we can use ${\mathcal{R}}(S)$ as a good measure for the conditioning of $S$. Note that the singular value truncation used in rank-adaptive methods ensures that $\varsigma_{r}(S)$ is always sufficiently large. Figure˜2a) and b) show the dynamics of ${\mathcal{R}}(S(t))$ and $\kappa(S(t))$ during low-rank regularized training; we see that $\kappa(S(t))$ decays as ${\mathcal{R}}(S(t))$ decays, validating ˜2.
Remark 1.
When $U,V$ are not orthonormal, e.g., in simultaneous gradient descent training (LoRA), Moreover, the smallest nrn-r singular values of $USV^{\top}$ are often zero-valued, and thus the bound of Equation˜4 is not useful. Table˜1 shows that the clean accuracy and adversarial accuracy of regularized LoRA is significantly lower than standard or regularized training with orthonormal $U,V$.
(a) $\kappa(S(t))$
(b) $\mathcal{R}(S(t))$
(c) $\widetilde{\mathcal{L}}(t)$
Figure 2: UCM Dataset, $\kappa(S(t))$ and $\mathcal{R}(S(t))$ of layer 4 of VGG16 for different regularizations strengths $\beta$. Each line is the median of 5 training runs. Higher $\beta$ values lead to faster reduction of the layer condition $\kappa(S)$, which quickly approaches its minimum value 11, and faster decay of $\mathcal{R}$. Unregularized training ($\beta=0$) leads to $\kappa(S)>1000$ after a few iterations.
We now study the stability of the regularizer when applied to a least squares regression problem, i.e., given a fixed $M\in\mathbb{R}^{r\times r}$ we seek to minimize $\mathcal{J}(S):=\beta{\mathcal{R}}(S)+\frac{1}{2}\|S-M\|^{2}$ over $S\in\mathbb{R}^{r\times r}$.
@@
Consider the dynamical system generated by the gradient flow of $\mathcal{J}$; namely, $\dot{S}(t)+\beta{\nabla}{\mathcal{R}}(S(t))+S(t)=M$. Then for any $t\geq 0$ we have the long-time stability estimate
$\tfrac{1}{2}\|S(t)-M\|^{2}+2\beta\int_{0}^{t}e^{\tau-t}{\mathcal{R}}(S(\tau))\,\mathrm{d}\tau\leq\tfrac{1}{2}e^{-t}\|S(0)-M\|^{2}+2(1-e^{-t})\beta(1+2\beta)\|M\|^{2}.$ (5)
See Appendix˜C for the proof. We note that unlike standard ridge and lasso regularizations methods, $\mathcal{R}$} lacks convexity; thus long-time stability of the regularized dynamics is not obvious. However, $\nabla$$\mathcal{R}$} possesses monotonicity properties that we leverage to show in (5) that the growth in $\mathcal{J}$ only depends on $\beta$, $M$, and the initial loss. Moreover, for large $t$, the change in the final loss by the regularizer only depends on $\beta$ and the true solution $M$ and not the specific path $S(t)$. While training on the non-convex loss will not provide the same theoretical properties as the convex least-square loss used in ˜3, the experiments in Figure˜2 give confidence that adding our regularizer does not yield a relatively large change in the loss decay rate over moderate training regimes. Particularly, we observe empirically in Figure˜2 that the condition number $\kappa(S)$ of decreases alongside the regularizer value $\mathcal{R}$ during training.